Find the total surface area of a cone with a radius of 5 m and a slant height of 8 m. Round to two decimal places.
Answer by Alan3354(69443)

Find the total surface area of a cone with a radius of 5 m and a slant height of 8 m. Round to two decimal places.
---------
Lateral area = pi*s*h = 40pi
--
Base area = pi*r^2